A particle doing simple harmonic motion whose amplitude is 4cm,time period is 12 sec.the ratio between time taken by it in going from its mean position to 2 cm and 2 cm to extreme position is

Grade:11

1 Answers

Arun
6 years ago
y = 4 sin(2 pi/12)*t
Now y = 2
2 = 4 sin(pi/6) *t1
1/2 = sin(pi/6) *t1
Pi/6 = (pi/6)* t1
t1 = 1 sec
Hence T1 = t1 - 0 = 1 sec
 
Now, similarily
Put y = 4
You will get t2 = 3 sec.
Hence T2 = t2 - t1 = 3- 1 = 2 sec
Hence
T1 : T2 = 1:2